# Dynamics-Aligned Latent Imagination in Contextual World Models for Zero-Shot Generalization

### v1

@@ 3 Preliminaries
To investigate zero-shot generalization in partially observable environments, we adopt a contextual Markov Decision Process (cMDP) framework, following Hallak et al. (2015); Modi et al. (2018); Kirk et al. (2023). A cMDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{C},\mathcal{R},\mathcal{P},\mathcal{E},\mu,p_{\mathcal{C}})$, where $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{O}$ are resp. the state, action, and observation spaces, $\mathcal{C}\subseteq\mathbb{R}^{d}$ is the context space with contexts $c\in\mathcal{C}$ drawn i.i.d. from a distribution $p_{\mathcal{C}}$, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{C}\to\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{C}\to\Delta(\mathcal{S})$ is the stochastic transition function specifying the probability distribution over next states given the current state, action, and context, $\mathcal{E}:\mathcal{S}\times\mathcal{C}\to\Delta(\mathcal{O})$ is the observation function specifying the distribution over observations given the state and context, $\mu:\mathcal{C}\to\Delta(\mathcal{S})$ is the initial state distribution where $\mu(s_{0}|c)$ specifies the probability of initial state $s_{0}$ given context $c$. In this partially observable setting, the agent does not observe the state $s_{t}$ or context $c$ directly but receives observations $o_{t}\in\mathcal{O}$. The objective is to learn a policy $\pi(a_{t}|o_{1:t},a_{1:t-1})$ that maximizes the expected return $\mathbb{E}\left[\sum_{t=0}^{T}r_{t}\right]$, where $T$ is the episode length.
@@ 4.3 Integrating Context into DreamerV3
Training. Training the context encoder loss (3) aligns ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}=g_{\varphi}(o_{t-K:t},a_{t-K:t-1})$ with environmental dynamics and requires careful handling of the DreamerV3 world model’s recurrence. Both integration strategies unroll the world model over a $K$-length window, initializing $h_{t-K}=\mathbf{0}$ (or a learned initial state), generating ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau}=g_{\varphi}(o_{\tau-K:\tau},a_{\tau-K:\tau-1})$ for $\tau=t-K$ to $t$. In Shallow Integration, $z_{\tau}\sim q_{\theta}(z_{\tau}|h_{\tau},o_{\tau},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau})$ and $h_{\tau+1}=f_{\theta}(h_{\tau},z_{\tau},a_{\tau})$; in Deep Integration, $z_{\tau}\sim q_{\theta}(z_{\tau}|h_{\tau},o_{\tau})$ and $h_{\tau+1}=f_{\theta}(h_{\tau},z_{\tau},a_{\tau},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau})$. For both, gradients through $h_{\tau}$ and $z_{\tau}$ are stopped in the recurrent dynamics, and through $h_{\tau}$ in the world model’s encoder, preventing updates to $\theta$. In Shallow Integration, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau}$ gradients are preserved in the world model’s encoder and losses $L_{\text{FD}}$ and $L_{\text{cross}}$, allowing updates to $\varphi$, $W_{z}$, and $W_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}}$. In Deep Integration, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau}$ gradients are stopped in the recurrent dynamics and preserved only in $L_{\text{FD}}$ and $L_{\text{cross}}$. This decouples context learning from the world model’s recurrence, ensuring ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}$ captures relevant dynamics for both strategies. For detailed pseudocode, see Algorithms 1 and 2 in Appendix B for Shallow Integration, and Algorithms 3 and 4 for Deep Integration.
@@ 2 Model components
     // Dynamics learning (World Model $\theta$)
@@ 18     Draw $B$ data sequences $\{(o_{t},a_{t})\}_{t=1}^{K+1}\sim\mathcal{D}$
1 Initialize dataset $\mathcal{D}$ with $S$ random seed episodes
@@
     // Dynamics learning (World Model $\theta$)
@@ 2 Model components
     // Dynamics learning (World Model $\theta$)
@@ 18     Draw $B$ data sequences $\{(o_{t},a_{t})\}_{t=1}^{K+1}\sim\mathcal{D}$
1 Initialize dataset $\mathcal{D}$ with $S$ random seed episodes
@@
     // Dynamics learning (World Model $\theta$)
@@ C.3 Architectural details
Context Encoder. The context encoder employs a standard transformer encoder block (Vaswani et al., 2017) to process a sequence of $K$ transitions, $(o_{t-K:t},a_{t-K:t-1})$, and produce the context representation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}\in\mathbb{R}^{8}$. The input sequence is fed directly into a dense layer with 256 units, bypassing a traditional embedding layer. This is followed by layer normalization and a single-head self-attention block with a skip connection from the dense layer’s output. The attention output is combined with the skip connection via a residual connection, followed by a second layer normalization. A two-layer MLP, each layer with 256 units, processes this output, with another residual connection combining the MLP output with the second layer norm’s output. A final dense layer projects the 256-dimensional intermediate representation to ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}\in\mathbb{R}^{8}$. All hidden layers, including the MLP and attention head, use 256 units and SiLU activation functions, consistent with DreamerV3 (Hafner et al., 2025). No masking is applied in the self-attention mechanism, as the input sequence is complete, enabling full contextual processing to generate ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}$.

### v2

@@ 3 Preliminaries
To investigate zero-shot generalization in partially observable environments, we adopt a contextual Markov Decision Process (cMDP) framework, following Hallak et al. (2015); Kirk et al. (2023). A cMDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{C},\mathcal{R},\mathcal{P},\mathcal{E},\mu,p_{\mathcal{C}})$, where $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{O}$ are resp. the state, action, and observation spaces, $\mathcal{C}\subseteq\mathbb{R}^{d}$ is the context space with contexts $c\in\mathcal{C}$ drawn i.i.d. from a distribution $p_{\mathcal{C}}$, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{C}\to\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{C}\to\Delta(\mathcal{S})$ is the stochastic transition function specifying the probability distribution over next states given the current state, action, and context, $\mathcal{E}:\mathcal{S}\times\mathcal{C}\to\Delta(\mathcal{O})$ is the observation function specifying the distribution over observations given the state and context, $\mu:\mathcal{C}\to\Delta(\mathcal{S})$ is the initial state distribution where $\mu(s_{0}|c)$ specifies the probability of initial state $s_{0}$ given context $c$. In this partially observable setting, the agent does not observe the state $s_{t}$ or context $c$ directly but receives observations $o_{t}\in\mathcal{O}$. The objective is to learn a policy $\pi(a_{t}|o_{1:t},a_{1:t-1})$ that maximizes the expected return $\mathbb{E}\left[\sum_{t=0}^{T}r_{t}\right]$, where $T$ is the episode length.
@@ 4.3 Integrating Context into DreamerV3
Training. Training the context encoder, the loss (3) aligns ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}=g_{\varphi}(o_{t-K:t},a_{t-K:t-1})$ with environmental dynamics and requires careful handling of the DreamerV3 world model’s recurrence. Both integration strategies unroll the world model over a $K$-length window, initializing $h_{t-K}=\mathbf{0}$ (or a learned initial state), generating ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau}=g_{\varphi}(o_{\tau-K:\tau},a_{\tau-K:\tau-1})$ for $\tau=t-K$ to $t$. In Shallow Integration, $z_{\tau}\sim q_{\theta}(z_{\tau}|h_{\tau},o_{\tau},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau})$ and $h_{\tau+1}=f_{\theta}(h_{\tau},z_{\tau},a_{\tau})$; in Deep Integration, $z_{\tau}\sim q_{\theta}(z_{\tau}|h_{\tau},o_{\tau})$ and $h_{\tau+1}=f_{\theta}(h_{\tau},z_{\tau},a_{\tau},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau})$. For both, gradients through $h_{\tau}$ and $z_{\tau}$ are stopped in the recurrent dynamics, and through $h_{\tau}$ in the world model’s encoder, preventing updates to $\theta$. In Shallow Integration, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau}$ gradients are preserved in the world model’s encoder and losses $L_{\text{FD}}$ and $L_{\text{cross}}$, allowing updates to $\varphi$, $W_{z}$, and $W_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}}$. In Deep Integration, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{\tau}$ gradients are stopped in the recurrent dynamics and preserved only in $L_{\text{FD}}$ and $L_{\text{cross}}$. This decouples context learning from the world model’s recurrence, ensuring ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}$ captures relevant dynamics for both strategies. For detailed pseudocode, see Algorithms 1 and 2 in Appendix B for Shallow Integration, and Algorithms 3 and 4 for Deep Integration.
@@ 7 Discussion and Outlook
Acknowledgments and Disclosure of Funding
The authors thank the anonymous reviewers for their valuable feedback and the NeurIPS community for fostering open and rigorous scientific exchange. JB and ME gratefully acknowledge funding by the German Research Foundation DFG through the MoReSpace (402776968) project.
@@ 2 Model components
7
    1ex// Dynamics learning (World Model $\theta$)
@@ 19     Draw $B$ data sequences $\{(o_{t},a_{t})\}_{t=1}^{K+1}\sim\mathcal{D}$
1Initialize dataset $\mathcal{D}$ with $S$ random seed episodes
@@
6
    1ex// Dynamics learning (World Model $\theta$)
@@ 2 Model components
7
    1ex// Dynamics learning (World Model $\theta$)
@@ 19     Draw $B$ data sequences $\{(o_{t},a_{t})\}_{t=1}^{K+1}\sim\mathcal{D}$
1Initialize dataset $\mathcal{D}$ with $S$ random seed episodes
@@
6
    1ex// Dynamics learning (World Model $\theta$)
@@ C.3 Architectural details
Figure 3 shows a high-level illustration of the DALI architecture.
Figure 3: DALI architecture overview.
Context Encoder. The context encoder $g_{\phi}$ employs a standard transformer encoder block (Vaswani et al., 2017) to process a sequence of $K$ transitions, $(o_{t-K:t},a_{t-K:t-1})$, and produce the context representation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}\in\mathbb{R}^{8}$. The input sequence is fed directly into a dense layer with 256 units, bypassing a traditional embedding layer. This is followed by layer normalization and a single-head self-attention block with a skip connection from the dense layer’s output. The attention output is combined with the skip connection via a residual connection, followed by a second layer normalization. A two-layer MLP, each layer with 256 units, processes this output, with another residual connection combining the MLP output with the second layer norm’s output. A final dense layer projects the 256-dimensional intermediate representation to ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}\in\mathbb{R}^{8}$. All hidden layers, including the MLP and attention head, use 256 units and SiLU activation functions, consistent with DreamerV3 (Hafner et al., 2025). No masking is applied in the self-attention mechanism, as the input sequence is complete, enabling full contextual processing to generate ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathfrak{z}}_{t}$.
@@ C.4 Computational Setup and Resource Requirements
C.5 Additional Experiments
Figure 4 shows the learning curves for DMC Ball-in-Cup and Walker Walk tasks under Featurized and Pixel-based observation modalities.
DALI’s context encoder adds only about $4\%$ more parameters to DreamerV3-small while scaling effectively to high-dimensional tasks. We evaluate it on DMC Quadruped Walk (56D observations, 12D actions), a larger environment than Walker Walk (24D, 6D). The 56D observation space emphasizes the challenge of torque-sensitive locomotion, where increased coordination and a higher number of actuators (12 vs. 6) demand more robust context inference. Trained for 600K timesteps and contextualized with gravity and actuator strength (same ranges as Walker Walk), DALI attains Featurized Extrapolation IQMs of $0.326\pm 0.043$ (DALI-S), and $0.389\pm 0.027$ (DALI-D-$\chi$), yielding up to $76.8\%$ improvement over the context-unaware baseline $0.220\pm 0.023$ (Dreamer-DR), and up to $50.8\%$ improvement over the context-aware baselines ($0.258\pm 0.028$, cRSSM-D; $0.317\pm 0.031$, cRSSM-S).
Ball-in-Cup
(a) Featurized
(b) Pixel
(c) Featurized
(d) Pixel
Walker Walk
Figure 4: Learning curves for DMC Ball-in-Cup and Walker Walk tasks under Featurized and Pixel-based observation modalities. Results show mean episode returns with $25$th–$75$th percentile confidence intervals.
@@ D.2 Actuator-Driven Recovery and Locomotion in Latent Imagination
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims presented in the abstract and introduction are fully supported by our theoretical and empirical findings, including a comprehensive analysis.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Limitations are discussed in the corresponding section of this paper.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: We provide a comprehensive proof of the necessity and benefit of our approach in Section 5 and Appendix A.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: This paper provides the complete code and instructions for running the experiments in a reproducible manner. Python dependencies, their versions, and all random seeds used are clearly specified.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: The full source code is provided with this submission along with all relevant details required to reproduce the presented experiments.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Yes, we provide all details for training the agent in-distribution and testing zero-shot generalization.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: For this purpose, we follow established methods for measuring the agent’s performance (Agarwal et al., 2021).
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: DALI adds only a small context encoder module to DreamerV3. As reported by Hafner et al. (2025), DreamerV3 can be trained on a single GPU, and our extension has the same requirement. Each trial with a single random seed takes roughly 1–2 days, depending on the observation modality.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: No disagreement amoung the team.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. Justification: [N/A]
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: [N/A]
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: This work builds upon and extends the codebase of Prasanna et al. (2024), which is properly credited. All associated licenses and terms of use from the original work are respected.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. 14.
    Crowdsourcing and research with human subjects
  53. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  54. Answer: [N/A]
  55. Justification: [N/A]
  56. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  57. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  58. Answer: [N/A]
  59. Justification: [N/A]
  60. 16.
    Declaration of LLM usage
  61. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  62. Answer: [N/A]
  63. Justification: [N/A]